OpenStudy (anonymous):

it suppose to be 0.7x+8≤ 1.1x-1

OpenStudy (angela210793):

1.7x-1.1x<=-1-8 0.6x<=-9 x<=-9/0.6

OpenStudy (anonymous):

1.7x-1.1x<=-1-8 0.6x<=-9 x<=-9/0.6 angela is right